(1+i)^100 how to solve without calculator
1)1
2)0
3)2496

Answer by Alan3354(69443)

(1+i)^100 how to solve without calculator
--------
(1+i)^2 = 1 + 2i - 1 = 2i
(1+i)^4 = (2i)^2 = -4
(1+i)^100 = (-4)^25